im trying to solve for change in thickness

the formula in words is;

compressibility= - (change in thickness/origional thickness)/ change in vertical effective stress.

I have everything but change in thickness, but i need to convert to the same units. could use a little help.

hers what I have:

compressibility= 6x10^-8 M^2/N

Origional thickness: 45 M (135 ft)

change in vertical effective stress= lb/ft^2


Thanks for the help
Answer #1

1lb = 0.453 Kg

1ft = 1/3.28 m

change in vertical effective stress= lb/ft^2 = 0.453x(3.28)^2= 4.873 Kg/m^2

as 1N = 1 Kg*m/s^2 , all units are of same type (Compressibility =6x10^-8 m^2/N, Original thickness = 45m , change in vertical effective stress= 4.873 Kg/m^2).

                                                                           OR                                                                          

1m = 3.28084 ft

1N = 0.224 808 943 87 pound-force = 0.224 lbF

Compressibility = 6x10^-8 x (3.280 ft)^2/(0.224 lbF) = 288.171 ft^2/lbF

Now all are same units.

Compressibility = 288.171 ft^2/lbF , Original thickness = 135 ft , change in vertical effective stress= lb/ft^2.